# Near-inertial echoes of ageostrophic instability in submesoscale filaments

### v2

@@ Abstract
Ocean submesoscales, flows with characteristic size around $10~{}\text{m}-10~{}\text{km}$, are transitional between the larger, rotationally-constrained mesoscale and three-dimensional turbulence. In this paper we present simulations of a submesoscale ocean filament. In our case, the filament is strongly sheared in both vertical and cross-filament directions and is unstable. Instability indeed dominates the early behaviour with a fast extraction of kinetic energy from the vertically sheared thermal wind. The filament is expected to be subject to instabilities, including symmetric and Kelvin-Helmholtz instabilities. However, the instability that emerges does not exhibit characteristics that match either, and appears to be non-normal in nature. The action of the instability is sufficiently rapid — with energy extraction from the mean flow emerging and peaking within the first inertial period ($\sim 18~{}\text{hr}$) — that the filament does not respond in a geostrophically balanced sense. Instead, it later exhibits vertically sheared near-inertial oscillations with higher amplitude as the initial minimum Richardson number decreases and at all initial noise levels. Horizontal gradients strengthen only briefly as the fronts restratify. These unstable filaments can be generated by strong mixing events at pre-existing stable structures; we also caution against inadvertently triggering this response in idealised studies that start in a very unstable state.
@@ 1 Introduction
Images of ocean properties at submesoscale resolution are often peppered with highly anisotropic features. Fronts — confined lateral gradients in density — and filaments — opposing pairs of fronts — remain balanced by the Coriolis acceleration of an along-front jet. In the open ocean, these features are created primarily at the edges of mesoscale vortices such as those created by separating western boundary currents, such as the Gulf Stream or Kuroshio current (McWilliams, 2016). As they are drawn out by the strain flow in these regions, horizontal gradients in velocities and tracers become larger, eventually leading to the onset of a new dynamical regime, where geostrophic and hydrostatic balances no longer describe the flow well (Shakespeare, 2016; Barkan et al., 2019; Taylor & Thompson, 2022).
In addition to the along-front geostrophic jet, an ageostrophic secondary circulation may develop at a front due to a background strain, turbulent mixing, instability or otherwise (Sullivan & McWilliams, 2018; Crowe & Taylor, 2018; McWilliams, 2021). The resulting vertical velocities provide a route for transport of tracers, momentum and energy between the surface layer and deep water (Mahadevan, 2016; Klein et al., 2019; Verma et al., 2019; Freilich & Mahadevan, 2021). Particularly, the submesoscale is a link between the larger scale, horizontally non-divergent flows and smaller scale turbulence that facilitates the observed forward cascade of energy in the ocean (Srinivasan et al., 2023).
In the absence of turbulence, horizontal gradients will approach infinity in finite time when forced by a strain flow (Hoskins & Bretherton, 1972). For a realistic front, some effective horizontal diffusion due to turbulence will arrest the frontogenesis before this singularity is reached, though this process is not well-understood (Samelson & Skyllingstad, 2016; Verma et al., 2019). The turbulence, namely the vertical momentum mixing induced by surface cooling or wind stress, acts on the vertically sheared thermal wind velocity. If only these processes act in the submesoscale mixed layer, the expected momentum balance is the turbulent thermal wind balance between the Coriolis acceleration of an along-front jet, the horizontal pressure gradient and vertical momentum mixing (Gula et al., 2014; Crowe & Taylor, 2018; Bodner et al., 2023).
@@
The behaviour of fluid susceptible to SI (or close to unstable) is suspected to have important consequences for ocean flows at the submesoscale and, for example, has been studied in the contexts of frontogenesis and frontal arrest (Thomas, 2012; Verma et al., 2019), mixing and vertical transport (Brannigan, 2016; Bosse et al., 2021), and near-inertial wave generation (Wienkers et al., 2021). However, the initial and long-time behaviour of a symmetrically unstable flow differ in form, growth rate and energy source (Heifetz & Farrell, 2007; Zemskova et al., 2020; Kimura, 2024). Initial motions in an unstable flow seeded with noise do not resemble the symmetric normal modes, and exhibit a faster growth rate across parameter space, a behaviour that is general across hydrodynamic stability (Trefethen et al., 1993). This non-normal, transient growth will play a particularly important role in this study.
In this paper, we present a set of strongly baroclinic ($Ri\leq 0.2$) simulations resembling the configuration of Sullivan & McWilliams (2018), but with coarser spatial resolution and without the surface cooling that drives vertical mixing. Our results focus on the spin-down of the fronts in the presence of noise and their subsequent instability. We show that instabilities act quickly to remove kinetic energy and excite near-inertial oscillations within the fronts that briefly strengthen horizontal buoyancy gradients at the surface. We argue that we observe a transient response of the filament associated with its symmetric instability. Finally, we suggest that initial conditions of numerical studies of similar strongly baroclinic features must be designed with initial instability and possible subsequent oscillations in mind, or risk drawing conclusions that may not be generalisable to balanced flows in the real ocean.
@@ 2 Numerical simulations
The reference state is invariant in the down-front direction $y$ and the initial evolution introduces little variation along this direction for the large-scale velocity and buoyancy fields (i.e. the filament does not meander). As such, we define the mean state to simply be the down-front average of the full 3D state, with averaging operator
@@ 3.1 General evolution
which is half the downfront length of our domain. The small domain limits the baroclinic growth rate. Nevertheless, we ran a simulation with a longer down-front domain that resolved around eight wavelengths of this instability (not shown), in which the meandering indeed becomes noticeable sooner, at around $ft/2\pi\approx 2$. Nevertheless, we verified that even then, the baroclinic instability does not alter the conclusions we present in this study. The rest of this discussion is concerned with the evolution of the filament before the meandering, which can be understood from the perspective of the down-front mean state.
@@ 3.2 Initial evolution is a fast instability
Figure 6: At $t/2\pi=0.30$, comparison of an $x$-$z$ slice at $y=0$ of a) vertical velocity $w$ and b) contribution to vertical shear production in the left side of the filament for the simulation with $Ri_{\text{min}}=0.1$. The red, dashed line in the top left panel locates $z=-50\text{m}$, the depth of the $x$-$y$ slice of vertical velocity displayed in c). The horizontal slice of $w$ is coarse-grained with a Gaussian filter with a half-width of one grid cell for clarity. A video timeseries of this plot appears in the supplementary material as Movie 1.
@@ 3.3 The initial instability echoes throughout the front as an inertial oscillation
The secondary circulation is correlated with the surface front strength. The maximum front strength in figure 4a emerges at $ft/2\pi\approx 1$, at the same time that isopycnals begin to move apart. This coincides with the moment $\frac{\text{d}^{2}C}{\text{d}t^{2}}=0$, and the circulation changes sign (figure 8a).
@@ 4 Discussion and conclusions
The specific nature of the instability is outside the scope of this paper, but will no doubt be of interest to readers. In short, we believe it to be a non-normal, transient response towards SI that is sensitive to the initial noise, rather than normal-mode SI. While we leave a full demonstration to future studies, we find a comparison with reduced-noise simulations particularly illuminating. We present these results in appendix B. The initial noise illustrated in figure 2 has a large magnitude, compared to the mean flow; and its vertical profile is responsible for the apparent downward propagation of the VSP contribution from figures 6 to 7. At low levels of initial noise, the unstable modes that emerge are down-front symmetric (figure 9) before being destroyed by secondary instabilities (not shown). Perturbations emerge at the same rate throughout the mixed layer in a normal-mode fashion, rather than growing fastest at the surface. The low-noise behaviour is similar to the uniform-gradient Eady state studied by Wienkers et al. (2021). The authors observe a similar instability-to-inertial-oscillations pathway that begins with a strong symmetric instability in 2.5D direct numerical simulations of finite-width, zero Richardson number frontal zones (i.e., constant lateral buoyancy gradients in domains that are periodic in the cross-front direction).
Arobone & Sarkar (2015), Zemskova et al. (2020), and Kimura (2024) investigate the initial and long-time behaviour of a symmetrically unstable flow. The initial and normal growth of perturbations in such a flow are different. In particular, the perturbations are initially strongly asymmetric and not aligned with isopycnals, before becoming nearly symmetric for large times (Arobone & Sarkar, 2015). The instability at the start of the simulations in this paper is such an initial response of the filament associated with SI and, due to the large initial noise, the perturbations do not have time to grow into normal-mode SI before being isotropised by secondary instabilities acting on the growing modes, leading to the rapid energy extraction from the mean state via the VSP.
@@ Appendix A Parametrisation of the reference state
and mixed layer transition function $g(\xi)=\ln(1+e^{\xi})$. $N_{0}$ and $N$ are deep-water and mixed layer buoyancy frequencies respectively, $f$ is the Coriolis frequency, $L$ is the distance between front centres, $H$ is the depth of the mixed layer and $\ell$ is the frontal half-width. We parametrise the state by varying $N$ and $N_{0}$ to achieve prescribed minimum local Rossby and Richardson numbers, given respectively by
$Ro_{\text{min}}=\frac{(N^{2}-N_{0}^{2})\beta^{2}}{f^{2}\alpha^{2}}\max_{s}\Big{(}\gamma\gamma_{s}\Big{)}\quad\text{and}\quad Ri_{\text{min}}=\frac{f^{2}\alpha^{2}N^{2}}{\beta^{2}(N^{2}-N_{0}^{2})^{2}}\min_{s}\Big{(}\gamma_{s}^{-2}\Big{)}.$ ($)13𝑎𝑏$
The full set of input parameters is shown in table 1.
@@ Appendix B Transient behaviour and the impact of initial noise
For the simulation with smallest pre-initial noise amplitude, the form of the vertical velocity at early times is shown in figure 9. Characteristic SI modes are recovered. These modes are somewhat tilted with respect to the isopycnals, expected for non-hydrostatic SI (Wienkers et al., 2021).
@@
Following the TKE production by $\text{VSP}_{0}$, inertial oscillations are also observed in the secondary circulations, as seen in figure 11. Hovmöller plots in figure 12 show that the near-surface isopycnals follow the inertial oscillations, with a large front strength and enhanced dissipation just after the filament reaches a minimum width ($ft/2\pi\approx 1.7$). The generation of inertial oscillations and subsequent surface convergence and frontal strengthening is agnostic to the method of extraction of kinetic energy, as long as it is rapid enough.

### v4

@@ Abstract
Ocean submesoscales, flows with characteristic size around $10~{}\text{m}-10~{}\text{km}$, are transitional between the larger, rotationally-constrained mesoscale and three-dimensional turbulence. In this paper we present simulations of a submesoscale ocean filament. In our case, the filament is strongly sheared in both vertical and cross-filament directions and is unstable. Instability indeed dominates the early behaviour with a fast extraction of kinetic energy from the vertically sheared thermal wind. However, the instability that emerges does not exhibit characteristics that match the perhaps expected symmetric or Kelvin-Helmholtz instabilities, and appears to be non-normal in nature. The prominence of the transient response depends on the initial noise and, for large initial noise amplitudes, saturates before SI normal modes are able to develop. The action of the instability is sufficiently rapid — with energy extraction from the mean flow emerging and peaking within the first inertial period ($\sim 18~{}\text{hr}$) — that the filament does not respond in a geostrophically balanced sense. Instead, at all initial noise levels, it later exhibits vertically sheared near-inertial oscillations with higher amplitude as the initial minimum Richardson number decreases. Horizontal gradients strengthen only briefly as the fronts restratify. These unstable filaments can be generated by strong mixing events at pre-existing stable structures; we also caution against inadvertently triggering this response in idealised studies that start in a very unstable state.
@@ 1 Introduction
Images of ocean properties at submesoscale resolution are often peppered with highly anisotropic features. Fronts — confined lateral gradients in density — and filaments — opposing pairs of fronts — remain balanced by the Coriolis acceleration of an along-front jet. In the open ocean, these features are created primarily at the edges of mesoscale vortices such as those created by separating western boundary currents, such as the Gulf Stream or Kuroshio current (McWilliams, 2016). As fronts and filaments are drawn out by the strain flow in these regions, horizontal gradients in velocities and tracers become larger, eventually leading to the onset of a new dynamical regime, where geostrophic and hydrostatic balances no longer describe the flow well (Shakespeare, 2016; Barkan et al., 2019; Taylor & Thompson, 2022).
In addition to the along-front geostrophic jet, an ageostrophic secondary circulation may develop at a front due to a background strain, turbulent mixing, instability or otherwise (Sullivan & McWilliams, 2018; Crowe & Taylor, 2018; McWilliams, 2021). The resulting vertical velocities provide a route for transport of tracers, momentum and energy between the surface layer and deep water (Mahadevan, 2016; Klein et al., 2019; Verma et al., 2019; Freilich & Mahadevan, 2021). In particular, the submesoscale is a link between the larger scale, horizontally non-divergent flows and smaller scale turbulence that facilitates the observed forward cascade of energy in the ocean (Srinivasan et al., 2023).
In the absence of turbulence, finite time singularities form in the horizontal gradients when forced by a strain flow (Hoskins & Bretherton, 1972). For a realistic front, some effective horizontal diffusion due to turbulence will arrest the frontogenesis before this singularity is reached, though this process is not well-understood (Samelson & Skyllingstad, 2016; Verma et al., 2019). The turbulence, namely the vertical momentum mixing induced by surface cooling or wind stress, acts on the vertically sheared thermal wind velocity. If only these processes act in the submesoscale mixed layer, the expected momentum balance is the turbulent thermal wind balance between the Coriolis acceleration of an along-front jet, the horizontal pressure gradient and vertical momentum mixing (Gula et al., 2014; Crowe & Taylor, 2018; Bodner et al., 2023).
@@
The behaviour of fluid susceptible to SI (or close to unstable) is suspected to have important consequences for ocean flows at the submesoscale and, for example, has been studied in the contexts of frontogenesis and frontal arrest (Thomas, 2012; Verma et al., 2019), mixing and vertical transport (Brannigan, 2016; Bosse et al., 2021), and near-inertial wave generation (Wienkers et al., 2021a). However, the initial and long-time behaviour of a symmetrically unstable flow differ in form, growth rate and energy source. The form of the linear perturbations that optimally extract energy from an Eady background state differ from the most unstable normal mode during initial evolution and their growth rate is enhanced during an initial, transient phase of instability. This behaviour persists in fully non-linear, three-dimensional simulations (Heifetz & Farrell, 2007; Zemskova et al., 2020; Kimura, 2024). Initial motions in an unstable flow seeded with noise do not resemble the symmetric normal modes, and exhibit a faster growth rate across parameter space, a behaviour that is general across hydrodynamic stability (Trefethen et al., 1993; Zemskova et al., 2020). This non-normal, transient growth will play a particularly important role in this study.
In this paper, we present a set of strongly baroclinic ($Ri\leq 0.2$) simulations resembling the configuration of Sullivan & McWilliams (2018), but with coarser spatial resolution and without the surface cooling that drives vertical mixing. On the other hand, we run more simulations, allowing for a better parameter space exploration, and we do so for longer periods of time, namely, several inertial periods each time.
As in the aforementioned study, the reference state is unstable; and we pay particular attention to the consequences of instability when initial noise is applied and how this may obfuscate conclusions that assume a balanced turbulent thermal wind state. Specifically, we share the goal of Sullivan & McWilliams (2018) to have realistic levels of noise, but while they attempt to generate realistic turbulent noise by continuously forcing the ocean surface, we simply generate random noise that aims to mirror the turbulent kinetic energy profile of mixed layer convection driven by surface cooling. This also results in a larger amplitude than many idealised studies, which seed the fluid with very small levels of noise (Zemskova et al., 2020; Wienkers et al., 2021b).
We show that instabilities act quickly to remove kinetic energy and excite near-inertial oscillations within the fronts that briefly strengthen horizontal buoyancy gradients at the surface. We argue that we observe a transient response of the filament as it evolves towards normal-mode symmetric instability. Finally, we suggest that the initialisation of both turbulence and an unstable reference state must be done with care and initial conditions of numerical studies of similar strongly baroclinic features must be designed with initial instability and possible subsequent oscillations in mind, or risk drawing conclusions that may not be generalisable to balanced flows in the real ocean.
@@ 2 Numerical simulations
The reference state is invariant in the down-front direction $y$ and the initial evolution introduces little variation along this direction for the large-scale velocity and buoyancy fields (i.e. the filament does not meander as the short down-front domain constrains the growth rate of baroclinic instabilities). As such, we define the mean state to simply be the down-front average of the full 3D state, with averaging operator
@@ 3.1 General evolution
which is half the downfront length of our domain. The small domain limits the baroclinic growth rate. Nevertheless, we ran a simulation with a longer down-front domain that resolved around eight wavelengths of this instability (not shown), in which the meandering indeed becomes noticeable sooner, at around $ft/2\pi\approx 2$; and verified that, even then, the baroclinic instability does not alter the conclusions we present in this study. The rest of this discussion is concerned with the evolution of the filament before the meandering, which can be understood from the perspective of the down-front mean state.
@@ 3.2 Initial evolution is a fast instability
Figure 6: At $t/2\pi=0.30$, comparison of an $x$-$z$ slice at $y=0$ of a) vertical velocity $w$ and b) contribution to vertical shear production in the left side of the filament for the simulation with $Ri_{\text{min}}=0.0$. The red, dashed line in the top left panel locates $z=-50\text{m}$, the depth of the $x$-$y$ slice of vertical velocity displayed in c). The horizontal slice of $w$ is coarse-grained with a Gaussian filter with a half-width of one grid cell for clarity. A video timeseries of this plot appears in the supplementary material as Movie 1.
@@ 3.3 The initial instability echoes throughout the front as an inertial oscillation
The secondary circulation is correlated with the surface front strength. The maximum front strength in figure 4a emerges at $ft/2\pi\approx 1$, at the same time that isopycnals begin to move apart. This coincides with the moment ${\text{d}^{2}C}/{\text{d}t^{2}}=0$, and the circulation changes sign (figure 8a).
@@ 4 Discussion and conclusions
The specific nature of the instability is outside the scope of this paper, but will no doubt be of interest to readers. In short, we believe it to be a non-normal, transient response towards SI that is sensitive to the initial noise, rather than normal-mode SI. While we leave a full demonstration to future studies, we find a comparison with reduced-noise simulations particularly illuminating. We present these results in appendix B. The initial noise illustrated in figure 2 has a large magnitude, compared to the mean flow; and its vertical profile is responsible for the apparent downward propagation of the VSP contribution from figures 6 to 7. At low levels of initial noise, the unstable modes that emerge are down-front symmetric (figure 9) before being destroyed by secondary instabilities (not shown). Perturbations emerge at the same rate throughout the mixed layer in a normal-mode fashion, rather than growing fastest at the surface. The low-noise behaviour is similar to the uniform-gradient Eady state studied by Wienkers et al. (2021a). The authors observe a similar instability-to-inertial-oscillations pathway that begins with a strong symmetric instability in 2.5D direct numerical simulations of finite-width, zero Richardson number frontal zones (i.e., constant lateral buoyancy gradients in domains that are periodic in the cross-front direction). The generation of inertial oscillations and subsequent surface convergence and frontal strengthening is agnostic to the method of extraction of kinetic energy, as long as it is rapid enough.
Arobone & Sarkar (2015), Zemskova et al. (2020), and Kimura (2024) investigate the initial and long-time behaviour of a symmetrically unstable flow. The initial and normal growth of linear perturbations in such a flow are different. In particular, the perturbations are initially strongly asymmetric and not aligned with isopycnals, before becoming nearly symmetric for large times (Arobone & Sarkar, 2015). The instability at the start of the simulations in this paper is such an initial response of the filament associated with SI. Non-normal perturbations, seeded by the initial noise, do not have time to grow into normal-mode SI before becoming significantly non-linear and being isotropised by secondary instabilities acting on the growing modes and by interaction with the remainder of the initial noise, leading to the rapid energy extraction from the mean state via the VSP. The role and applicability of linear instability analyses in the real ocean, where internal waves, inertial waves and convection contribute to noise, which may strongly interact with both normal and transient forms of instability, is not clear and requires attention in future studies.
@@ Appendix A Parametrisation of the reference state
and mixed layer transition function $g(\xi)=\ln(1+e^{\xi})$. $N_{0}$ and $N$ are deep-water and mixed layer buoyancy frequencies respectively, $f$ is the Coriolis frequency, $L$ is the distance between front centres, $H$ is the depth of the mixed layer and $\ell$ is the frontal half-width. Remaining, non-dimensional, parameters are described in table 1. We parametrise the state by varying $N$ and $N_{0}$ to achieve prescribed minimum local Rossby and Richardson numbers, given respectively by
$Ro_{\text{min}}=\frac{(N^{2}-N_{0}^{2})\beta^{2}}{f^{2}\alpha^{2}}\max_{s}\Big{(}\gamma\gamma_{,s}\Big{)}\quad\text{and}\quad Ri_{\text{min}}=\frac{f^{2}\alpha^{2}N^{2}}{\beta^{2}(N^{2}-N_{0}^{2})^{2}}\min_{s}\Big{(}\gamma_{,s}^{-2}\Big{)}.$ ($)13𝑎𝑏$
Target Rossby and Richardson numbers are given in table 1.
@@ Appendix B Transient behaviour and the impact of initial noise
For the simulation with smallest pre-initial noise amplitude, the form of the vertical velocity at early times is shown in figure 9, and a video timeseries of vertical velocity slices and VSP (as in figures 6 and 7) appears in the supplementary material as Movie 2. Characteristic SI modes are recovered. These modes are somewhat tilted with respect to the isopycnals in the x-z plane, expected for non-hydrostatic SI (Wienkers et al., 2021a).
@@
Following the TKE production by $\text{VSP}_{0}$, inertial oscillations are also observed in the secondary circulations, as seen in figure 11. Hovmöller plots in figure 12 show that the near-surface isopycnals follow the inertial oscillations, with a large front strength and enhanced dissipation just after the filament reaches a minimum width ($ft/2\pi\approx 1.7$).